If one charge attracts the other with a force of 0.140N, what are the magnitudes of the two charges if their total charge is also 21.0 μC? The charges are at a distance of 2.10 m apart. Note that you may need to solve a quadratic equation to reach your answer. Enter the charge with a smaller magnitude in the first box. Q14.05 uC Q2= 16.9 uC Submit Answer Incorrect. Tries 6/10 Previous Tries
Answer #1

Here ,

let the charges are x and y

x + y = 21 ----(1)

k * x * y/d^2 = force

9*10^9 * x * y*10^-12/2.1^2 = 0.140 ---(2)

solving for x and y

x = 4.05 uC

y = 16.95 uC

the charges are 4.05 *10^-6 C and 16.95 *10^-6 C